OpenStudy (anonymous):

What does x equal? x-7 over x-8 is = to 1 over x-8.

OpenStudy (chaise):

\[\frac{x-7}{x-8}=\frac{1}{x-8}\] The equation has no solutions.

OpenStudy (anonymous):

how do you work that out chaise?(cringe...) surely this reduces to x-7 = 1, therefore, x=8? No? I know I'm walkin on ice here.

OpenStudy (stormfire1):

BPDlkeme is right...it's completely solvable and the answer is x = 8.